0 0
Add a comment Improve this question Transcribed image text
Answer #1

a)
C= q/ΔV

C = (0.0153*10^-6C)/(200) = 0.765*10^-10F

b)
C = ϵo A/ d

A = Cb / ϵo

A = (0.765*10^-10)*(2.15*10^-3)/(8.854*10^-12C/Nm^2) = 0.0185m

c)

EST = 1/2 QΔV

EST = 1/2 (0.0153*10^-6)(200) = 1.53*10^-6j

d)

3e6 v/m x 0.0215 m = 64500 volts